OpenStudy (anonymous):

Write the rule for this table as an equation. s a 7 304 8 356 9 408 10 460

OpenStudy (anonymous):

seems to be a linear function.

OpenStudy (anonymous):

simple AP series

OpenStudy (anonymous):

is it a sequence or a series?

OpenStudy (anonymous):

sequence Dr Technical :P

OpenStudy (anonymous):

Pl ease help!

OpenStudy (anonymous):

first term = a common difference d = 52 so t= a + 52d-52

OpenStudy (anonymous):

or, t = 252 +52d-52 that is t =200=52d

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so then?

OpenStudy (anonymous):

is that the final answer?